OpenStudy (anonymous):

SEVEN+SEVEN+SIX=TWENTY Each letter represents a unique digit. "SEVEN" is a five digit number. Find a possible solution.

OpenStudy (lgbasallote):

E = 7

OpenStudy (lgbasallote):

S = 3

OpenStudy (anonymous):

@myininaya @TuringTest @FoolForMath Math brain teaser. :P @KingGeorge @Mr.Math

OpenStudy (anonymous):

More like crypto, but okay.

OpenStudy (lgbasallote):

^he just snobbed my answers -___-

OpenStudy (anonymous):

Your answers don't work. :P

OpenStudy (lgbasallote):

oh..darn..

OpenStudy (kinggeorge):

I don't have too much time here, but let's see. We know that N+N+X=?Y Where ? may be 1 or 0.

OpenStudy (kinggeorge):

E+E+I+?=$T Where $ could be 0, 1, or 2.

OpenStudy (kinggeorge):

At the other side, we have S+S=T% Where % is either W or W-1.

OpenStudy (kinggeorge):

So \(S\geq5\) And T=1.

OpenStudy (anonymous):

Correct so far. :D Although, I am interested in seeing the solution, as only the question and the answer are posted.

OpenStudy (lgbasallote):

oh darn.. s>= 5..im wrong lol

OpenStudy (kinggeorge):

Let's look at EN+EN+IX=%1Y Using the same percent from above. Maybe we can see something there.

OpenStudy (kinggeorge):

So either E+E+I is divisible by 10, or \(N+N+X\geq10\)

OpenStudy (kinggeorge):

I've got to go now, but good luck to every looking at this.

OpenStudy (anonymous):

You've been quite helpful in getting me to understand approaching this problem. :D

OpenStudy (anonymous):

Alphamatics or cryptogram.

OpenStudy (anonymous):

You know, I find these sorts of problems really cool. Just never ever had them before.

OpenStudy (anonymous):

One of my favorite is SLIDE- DEAN = 3651.

OpenStudy (anonymous):

I have a bucket full of these questions, with answers, but without solution manuals. :( I'm going through them, but they're haaard. Because I've never had any before.

OpenStudy (anonymous):

@Ishaan94 Might be interested?

OpenStudy (anonymous):

Use a program :P

OpenStudy (anonymous):

Hey, I already have the answers. >:|

OpenStudy (anonymous):

And WolframAlpha can't step by step this problem. :O

OpenStudy (anonymous):

I know but brute-forcing is fun ;)

OpenStudy (anonymous):

@KatrinaKaif If you know any crypto, it would be lovely. :D

OpenStudy (anonymous):

Whoops, wrong individual. @blues

OpenStudy (anonymous):

lolol what a mistake

OpenStudy (anonymous):

So, how do you solve this?